The greater of two numbers is one less than 8 times the smaller. Their sum is 98.
Answer by alicealc(293)

the smaller number = x
the greater number = 8x - 1

their sum is 98:
x + 8x - 1 = 98
9x - 1 = 98
9x = 98 + 1
9x = 99
x = 99/9 = 11

8x - 1 = 8*11 - 1 = 88 - 1 = 87

so, the numbers are 11 and 87